# Exact Random Graph Matching with Multiple Graphs

### v1

@@ 8                   Set $\widehat{\pi}_{1j}(v)=v^{\prime}$
Algorithm 2 runs in two steps: In step 1, the $k$-core estimator, for a suitable choice of $k$, is used to pairwise match all the graphs. For any $i$ and $j$, the $k$-core estimator selects a permutation $\widehat{\nu}_{ij}$ to maximize the size of the $k$-core [^1] The $k$-core of a graph $G$ is the largest subset of vertices $\mathsf{core}_{k}(G)$ such that the induced subgraph has minimum degree at least $k$. of $G_{i}\wedge G_{j}^{\widehat{\nu}_{ij}}$. It then outputs a matching $\widehat{\mu}_{ij}$ by restricting the domain of $\widehat{\nu}_{ij}$ to $\mathsf{core}_{k}(G_{i}\wedge G_{j}^{\widehat{\nu}_{ij}})$. These matchings $\widehat{\mu}_{ij}$ need not be complete - in fact, each of them is a partial matching with high probability whenever $Cs^{2}<1$. In step 2, these partial matchings are boosted as follows: If a node $v$ is unmatched between two graphs $G_{i}$ and $G_{j}$, then search for a sequence of graphs $G_{i},G_{k_{1}},\cdots,G_{k_{\ell}},G_{j}$ such that $v$ is matched between any two consecutive graphs in the sequence. If such a sequence exists, then extend $\widehat{\mu}_{i,j}$ to include $v$ by transitively matching it from $G_{i}$ to $G_{j}$.
In Section 4.2, we show that Algorithm 2 correctly matches all nodes across all graphs with probability $1-o(1)$, whenever the necessary condition $Cs(1-(1-s)^{m-1})>1$ holds. We remark that this also implies that Algorithm 1 succeeds under the same condition, because the MLE is optimal. Note that the MLE selects all permutations $\widehat{\pi}_{12},\cdots,\widehat{\pi}_{1m}$ simultaneously based on their union graph. In contrast, Algorithm 2 only ever makes pairwise comparisons between graphs. Perhaps surprisingly, it turns out that this is sufficient for exact recovery. An analysis of Algorithm 2 is presented in Section 4. Along the way, independent results of interest on the $k$-core of Erdős-Rényi graphs are obtained.
4 Proof Outlines and Key Insights
4.1 Impossibility of exact graph matching (Theorem 2)
This result has a simple proof following a genie-aided converse argument. The idea is to reduce the problem to that of matching two graphs by providing extra information to the estimator.
Proof of Theorem 2. If the correspondences $\pi^{*}_{12},\cdots,\pi^{*}_{1,m-1}$ were provided as extra information to an estimator, then the estimator must still match $G_{m}$ with the union graph $G_{1}^{\prime}\vee G_{2}^{\prime}\vee\cdots\vee G_{m-1}^{\prime}$. This can be viewed as an instance of matching two graphs obtained by asymmetric subsampling: the graph $G_{m}$ is obtained from a parent graph $G\sim\mathsf{ER}(n,C\log(n)/n)$ by subsampling each edge independently with probability $s_{1}:=s$, and the graph $\widetilde{G}_{m-1}:=G_{1}^{\prime}\vee G_{2}^{\prime}\vee\cdots\vee G_{m-1}^{\prime}$ is obtained from $G$ by subsampling each edge independently with probability $s_{2}:=1-(1-s)^{m-1}$. Cullina and Kiyavash studied this model for matching two graphs: Theorem 2 of [CK17] establishes that matching $G_{m}$ and $\widetilde{G}_{m-1}$ is impossible if $Cs_{1}s_{2}<1$, or equivalently if $Cs(1-(1-s)^{m-1})<1$. ∎
4.2 Achievability of exact graph matching (Theorem 3)

### v2

@@ 9                   Set $\widehat{\pi}_{1j}(v)=v^{\prime}$
  • Algorithm 1 only ever makes pairwise comparisons between graphs. In contrast, the maximum likelihood estimator for the simpler homogeneous setting selects the permutations $\widehat{\pi}_{12},\cdots,\widehat{\pi}_{1m}$ simultaneously based on all the graphs. Despite this, Algorithm 1 works down to the threshold for exact recovery.
  • The transitive closure sub-routine (step 2) has low computational complexity, $O(m^{2}n\log n)$. Further, it can be used to boost the pairwise matchings obtained by any estimator in step 1, i.e. it can be used as a black-box bridge between binary graph matching and $m$-ary graph matching whenever partial pairwise matchings are available. Our choice of the $k$-core estimator for step 1 is motivated by its precision, i.e. the $k$-core estimator does not make mistakes with high probability (CKMP, 19; RS, 23). This is crucial to avoid compounding errors during the boosting step. This precision is quantified formally later in the analysis (see Section 5.1.1).
4 Preliminaries: On the $k$-core of Inhomogeneous Graphs
This section presents useful preliminaries and lemmas for the proof of Theorem 1. The results are stated for general inhomogeneous graphs and may be of interest more generally beyond the scope of the graph alignment problem. The main objective of this section is to establish that under appropriate conditions, all nodes with sufficiently large degree in an inhomogeneous graph are in its $k$-core with high probability.
Lemma 13
Let $G\sim\mathrm{RG}(n,\mathbf{P})$ and $d_{i}=\sum_{j\neq i}p_{ij}$. Suppose there exist $\alpha\in(0,1]$ and $\varepsilon>0$ such that $p_{\max}=o(n^{\alpha/2-\varepsilon-1})$ and
$\sum_{i=1}^{n}e^{-d_{i}}=o(n^{-\alpha}).$
Then, for any positive integer $k$ and any $v\in[n]$,
($\mathbb{P}\left(\left\{v\notin\mathrm{core}_{k}(G)\right\}\cap\left\{\delta_{G}(v)\geq k+2/\alpha\right\}\right)=o(1/n).$
We build up to the proof of this lemma by establishing some results about the number of low-degree nodes (in Section 4.1), and about the size of the $k$-core in inhomogeneous random graphs (in Section 4.2). These results are then used to prove Lemma 13 in Section 4.3.
4.1 On low degree nodes in an inhomogeneous graph
Lemma 14
Let $p\in[0,1]$. Consider independent random variables $X$, $Y$ and $Z$ with distributions
$X\sim\mathrm{Bern}(p),\qquad Y\sim\mathrm{Pois}(p),\qquad Z\sim 2\cdot\mathrm{Pois}(p/2).$
Then, for all $t$, the moment generating functions of $X$, $Y$ and $Z$ are ordered as
$M_{X}(t)\leq M_{Y}(t)\leq M_{Z}(t).$
Proof. We have
$M_{X}(t)=1+p(e^{t}-1)\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\exp\left(p(e^{t}-1)\right)=M_{Y}(t)\stackrel{{\scriptstyle\text{(b)}}}{{\leq}}\exp\left(p\left(\frac{e^{2t}-1}{2}\right)\right)=M_{Z}(t),$
[…]
Lemma 15
Let $n>0$ and let $X_{1},\cdots,X_{n}$ be independent random variables with $X_{i}\sim\text{Bern}(p_{i})$. Let $m$ be an integer so that $0\leq m\leq n$, and define
$X:=\sum_{i\leq m}X_{i}+\sum_{j>m}2X_{j}.$
Then, for any $t<\mathbb{E}X$,
$\mathbb{P}\left(X\leq t\right)\leq\exp\left(-\frac{\mathbb{E}X}{2}+\frac{t}{2}\log\left(\frac{e\cdot\mathbb{E}X}{t}\right)\right)$
Proof. Let $Y_{i}\sim\text{Pois}(p_{i})$ and $Z_{i}\sim 2\cdot\text{Pois}(p_{i}/2)$. It follows that
$M_{X}(t)$ $=\prod_{i=1}^{m}M_{X_{i}}(t)\cdot\!\!\prod_{j=m+1}^{n}\!\!M_{X_{j}}(2t)$
$\leq\prod_{i=1}^{m}M_{Z_{i}}(t)\cdot\!\!\prod_{j=m+1}^{n}\!\!M_{Y_{j}}(2t)$
$=\prod_{i=1}^{m}M_{Z_{i}}(t)\cdot\!\!\prod_{j=m+1}^{n}\!\!M_{2Y_{j}}(t)=M_{Z}(t),$
where $Z:=\sum_{i=1}^{m}Z_{i}+\sum_{j=m+1}^{n}2Y_{j}$ is distributed as $Z\sim 2\cdot\text{Pois}\left(\mathbb{E}X/2\right)$.
Therefore, Chernoff tail bounds that hold for $Z$ also hold for $X$. In particular, for any $t<\mathbb{E}X$,
$\mathbb{P}\left(X\leq t\right)\leq\inf_{\theta\geq 0}e^{\theta t}M_{X}(-\theta)\leq\inf_{\theta\geq 0}e^{\theta t}M_{Z}(-\theta)=\inf_{\theta\geq 0}\exp\left(\theta t+\frac{\mathbb{E}{X}}{2}\left(e^{-2\theta}-1\right)\right).$
[…]
Lemma 16
Let $G\sim\mathrm{RG}(n,\mathbf{P})$, and let $d_{i}=\sum_{j\neq i}p_{ij}$. Let $r>0$ be an integer and let $Z_{r}$ denote the set of nodes of $G$ with degree less than or equal to $r$. Suppose that for some $\alpha>0$,
$\sum_{i=1}^{n}e^{-d_{i}}=o(n^{-\alpha}).$ (9)
  • For any $\delta>1-\alpha/2$,
    $\mathbb{P}\left(|Z_{r}|>n^{\delta}\right)=o(1/n).$ (10)
  • If $p_{\max}=o(n^{\alpha/2-1})$, then (10) is true for all $\delta>1-\alpha$.
Proof. If strictly more than $n^{\delta}$ nodes all have degree less than or equal to $r$, then a subset of exactly $n^{\delta}$ nodes has average degree less than $r$, i.e.
$\mathbb{P}\left(|Z_{r}|>n^{\delta}\right)$ $=\mathbb{P}\left(\exists\ T\subseteq[n]\colon\left\{|T|\geq n^{\delta}\right\}\cap\left\{\max_{u\in T}\delta_{G}(u)\leq r\right\}\right)$
$\leq\mathbb{P}\left(\exists\ T\subseteq[n]\colon\left\{|T|=n^{\delta}\right\}\cap\left\{\max_{u\in T}\delta_{G}(u)\leq r\right\}\right)$
$\leq\mathbb{P}\left(\exists\ S\subseteq[n]\colon\left\{|S|=n^{\delta}\right\}\cap\left\{\sum_{u\in S}\delta_{G}(u)\leq r|S|\right\}\right).$ (11)
Let $G_{uv}$ denote the indicator random variable for the presence of edge $\{u,v\}$ in $G$. For any set $S\subseteq[n]$, define
$\mu_{S}:=\sum_{i\in S}\mathbb{E}[\delta_{G}(i)]=\sum_{i\in S}d_{i}.$
Since for any $i\in[n]$, we have that $e^{-d_{i}}\leq\sum_{j=1}^{n}e^{-d_{j}}\leq n^{-\alpha},$ it follows that $d_{i}\geq\alpha\log n$. Therefore, for any $S\subseteq[n]$, we have that $\mu_{S}\geq\alpha|S|\log n$. Hence,
$\mathbb{P}\left(\sum_{u\in S}\delta_{G}(u)\leq r|S|\right)$ $\!=\mathbb{P}\left(\sum_{\begin{subarray}{c}u\in S\\ v\in[n]\setminus S\end{subarray}}\!\!\!G_{uv}+\!\!\sum_{\begin{subarray}{c}u,v\in S\\ u<v\end{subarray}}2G_{uv}\leq r|S|\right)\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\exp\left(\!-\frac{\mu_{S}}{2}\!+\!\frac{r|S|}{2}\log\left(\frac{e\cdot\mu_{S}}{r|S|}\right)\!\right)$
$\stackrel{{\scriptstyle\text{(b)}}}{{\leq}}\exp\left(-\frac{\alpha|S|}{2}\log(n)+\frac{r|S|}{2}\log\left(\frac{e\cdot\alpha\log n}{r}\right)\right)$ (12)
where (a) uses Lemma 15 and (b) uses that the function on the right hand side of (b) is decreasing in $\mu_{S}$ on the interval $(r|S|,\infty)$, and that $r|S|\leq\alpha|S|\log n\leq\mu_{S}$ for all sufficiently large $n$. Finally, combining (11) and (12), and using that $|S|=n^{\delta}$ yields
$\mathbb{P}\left(|Z_{r}|>n^{\delta}\right)$ $\leq\!\!\sum_{S\subseteq[n]:|S|=n^{\delta}}\!\!\!\!\mathbb{P}\left(\sum_{u\in S}\delta_{G}(v)\leq r|S|\right)\leq\binom{n}{n^{\delta}}\exp\left(-\frac{\alpha n^{\delta}}{2}\log n+\frac{rn^{\delta}}{2}\log\left(\frac{e\cdot\alpha\log n}{r}\right)\!\right)$
$\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\exp\left(\left(1-\delta-\frac{\alpha}{2}\right)n^{\delta}\log n+\frac{rn^{\delta}}{2}\log\left(\frac{e\cdot\alpha\log n}{r}\right)\right)=o(1/n),$
whenever $\delta>1-\alpha/2$. Here, (a) uses the fact that $\binom{n}{n^{\delta}}\leq\exp\left((1-\delta)n^{\delta}\log n\right)$. Hence, (10) follows.
Suppose that $p_{\max}=o(n^{\alpha/2-1})$. Notice that
$\sum_{u\in S}\delta_{G}(u)=\sum_{u\in S}\sum_{\begin{subarray}{c}v\in[n]\\ v\neq u\end{subarray}}G_{uv}\leq\sum_{u\in S}\sum_{\begin{subarray}{c}v\in[n]\setminus S\end{subarray}}G_{uv}\triangleq Y_{S},$
and so we have that
$\mathbb{P}\left(\sum_{u\in S}\delta_{G}(u)\leq r|S|\right)\leq\mathbb{P}\left(Y_{S}\leq r|S|\right)\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}e^{-\mathbb{E}\left[Y_{S}\right]}\cdot\left(\frac{e\cdot\mathbb{E}\left[Y_{S}\right]}{r|S|}\right)^{r|S|},$ (13)
where (a) uses that $Y_{S}$ is a sum of independent Bernoulli random variables, and so the Chernoff tail bound (19) can be applied. On the other hand,
$\mathbb{E}\left[Y_{S}\right]=\sum_{u\in S}d_{u}-\sum_{u\in S}\sum_{\begin{subarray}{c}v\in S\\ v\neq u\end{subarray}}2p_{uv}\stackrel{{\scriptstyle\text{(a)}}}{{\geq}}\sum_{u\in S}d_{u}-\binom{|S|}{2}\cdot(2p_{\max})\stackrel{{\scriptstyle\text{(b)}}}{{\geq}}\alpha|S|\log(n)-|S|^{2}\cdot p_{\max}\triangleq\widetilde{\mu}_{S},$
where (a) uses that $p_{uv}\leq p_{\max}$ and (b) uses that $\mu_{S}\geq\alpha|S|\log n$. Continuing from (13), we have
$\mathbb{P}\left(Y_{S}\leq r|S|\right)\leq e^{-\mathbb{E}\left[Y_{S}\right]}\cdot\left(\frac{e\cdot\mathbb{E}\left[Y_{S}\right]}{r|S|}\right)^{r|S|}\leq e^{-\widetilde{\mu}_{S}}\cdot\left(\frac{e\cdot\widetilde{\mu}_{S}}{r|S|}\right)^{r|S|},$ (14)
where the last step uses that the function $x\mapsto e^{-x}\cdot\left(\frac{e\cdot x}{r|S|}\right)^{r|S|}$ is decreasing in $x$ whenever $x\geq r|S|$. Thus, for any $S$ such that $|S|=n^{\delta}$, we have from (14) that
$\mathbb{P}\left(Y_{S}\leq rn^{\delta}\right)$ $\leq\exp\left(-n^{\delta}\left(\alpha\log(n)-r-n^{\delta}p_{\max}\right)\right)\cdot\left(\frac{\alpha\log(n)-n^{\delta}p_{\max}}{r}\right)^{rn^{\delta}}$
$=\exp\left(-n^{\delta}\left(\alpha\log(n)-r-n^{\delta}p_{\max}-r\log\left(\frac{\alpha\log(n)-n^{\delta}p_{\max}}{r}\right)\right)\right).$ (15)
Combining this with (11), a union bound yields that
$\mathbb{P}\left(|Z_{r}|>n^{\delta}\right)$ $\!\leq\mathbb{P}\left(\exists\ S\subseteq[n]:\left\{|S|=n^{\delta}\right\}\cap\left\{\sum_{u\in S}\delta_{G}(u)\leq r|S|\right\}\right)\leq\binom{n}{n^{\delta}}\mathbb{P}\left(Y_{S}\leq rn^{\delta}\right)$
$\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\!\exp\!\left(-n^{\delta}\Big{[}(\alpha\!+\!\delta\!-\!1)\log(n)-n^{\delta}p_{\max}-r\log\Big{(}\frac{\alpha\log(n)\!-\!n^{\delta}p_{\max}}{r}\Big{)}-r-1\Big{]}\right),$ (16)
[…]
4.2 On the size of the $k$-core of an inhomogeneous random graph
The following concentration inequalities for a sum of Bernoulli random variables are used frequently in the analysis.
Lemma 17
Let $X$ be a sum of $n$ independent Bernoulli random variables, and let $\mathbb{E}[X]=np$. Then,
  1. 1.
    For any $\delta>0$,
  2. 2.
    For any $\delta>5$,
  3. 3.
    For any $\delta\in(0,1)$,
    $\mathbb{P}\left(X\leq(1-\delta)np\right)\leq\left(\frac{e^{-\delta}}{(1-\delta)^{1-\delta}}\right)^{np}=e^{-np}\cdot\left(\frac{e}{1-\delta}\right)^{(1-\delta)np}.$ (19)
[…]
Let $G\sim\mathrm{RG}(n,\mathbf{P})$ and let $k>0$. Let $v$ be any node of $G$ and consider the induced subgraph $G-v$ of $G$ on the vertex set $[n]\setminus\{v\}$. The objective of this section is to show that the $k$-core of $G-v$ is sufficiently large with probability $1-o(1/n)$. Formally, we wish to prove the following.
Lemma 18
Let $G\sim\mathrm{RG}(n,\mathbf{P})$, and let $v$ be a node of $G$. Suppose there exist $\alpha\in(0,1]$ and $\varepsilon>0$ such that $p_{\max}=o(n^{\alpha/2-\varepsilon-1})$ and
$\sum_{i=1}^{n}e^{-d_{i}}=o(n^{-\alpha}).$
Then for any $\delta>1-\alpha$, the size of the $k$-core of $G-v$ satisfies
$\mathbb{P}\left(\left|\mathrm{core}_{k}(G-v)\right|<n-3n^{\delta}\right)=o(1/n).$
require :  Graph $G$, Set $U\subseteq V(G)$.
1 $U_{0}\leftarrow U$
2 for $i=0,1,2,3,\cdots$ do
4             $U_{i+1}\leftarrow U_{i}\cup\{u\}$
5
6       else
7             return $U_{i}$
8
9       end if
10
11 end for
Algorithm 2 Łuczak expansion
To this end, consider Algorithm 2 to identify a subset of the $k$-core, originally proposed by Łuczak (Łuc, 91). Note that the for loop eventually terminates - the set $V\setminus U_{i}$ is empty, for example, when $i=n$ for any input set $U$. The key is to realize that the for loop terminates quickly when the input $U$ is $Z_{k+1}$ , i.e. the set of vertices of the input graph $G$ whose degree is $k+1$ or less. Furthermore, for that input, the complement of the set output by the algorithm is contained in the $k$-core. This is formally stated in Lemma 19 in what follows.
Lemma 19
Let $G\sim\mathrm{RG}(n,\mathbf{P})$ and let $v$ be a node of $G$. Let $U_{f}$ be the output of Algorithm 2 with input graph $G-v$ and $U=Z_{k+1}$. Then,
  1. (a)
    $U_{f}^{c}\subseteq\mathrm{core}_{k}(G-v)$.
  2. (b)
    If there exist $\alpha\in(0,1]$ and $\varepsilon>0$ such that $p_{\max}=o(n^{\alpha/2-\varepsilon-1})$ and $\sum_{i=1}^{n}e^{-d_{i}}=o(n^{-\alpha})$, then for any $\delta>1-\alpha$,
    $\mathbb{P}\left(|U_{f}|>3n^{\delta}\right)=o(1/n).$
Proof. (a) The proof is by construction: Since $U_{f}$ is obtained by adding exactly $f$ nodes to $U_{0}$, it follows that $U_{f}^{c}\subseteq U_{0}^{c}=Z_{k+1}^{c}$, so each node in $U_{f}^{c}$ has degree $k+2$ or more in $G-v$. Further, each node in $U_{f}^{c}$ has at most $2$ neighbors in $U_{f}$, else the for loop would not have terminated. Thus, the subgraph of $G-v$ induced on the set $U_{f}^{c}$ has minimum degree at least $k$, and the result follows.
(b) Let $\delta^{\prime}$ be such that $1-\alpha<\delta^{\prime}<\min\left\{\delta,1-\alpha+2\varepsilon\right\}$. If $|U_{f}|>3n^{\delta}$, then either $|U_{0}|>3n^{\delta^{\prime}}$ or there is some $M$ in $\{0,1,\cdots,f\}$ for which $|U_{M}|=3n^{\delta^{\prime}}$. Therefore,
$\mathbb{P}\left(|U_{f}|>3n^{\delta}\right)$ $\leq\mathbb{P}\left(|U_{0}|>3n^{\delta^{\prime}}\right)+\mathbb{P}\left(\exists\ M\in\{0,1,\cdots,f\}\text{ s.t. }|U_{M}|=3n^{\delta^{\prime}}\right)$
$=o(1/n)+\underbrace{\mathbb{P}\left(\exists\ M\in\{0,1,\cdots,f\}\text{ s.t. }|U_{M}|=3n^{\delta^{\prime}}\right)}_{(\star)},$
where the first term is $o(1/n)$ by Lemma 16. Note that each iteration $i=0,1,\cdots,M-1$ of the for loop adds exactly $1$ vertex and at least $3$ edges to the subgraph of $G-v$ induced on $U_{M}$. Therefore, the induced subgraph $G|_{U_{M}}$ has $3n^{\delta}$ vertices and at least $3\left(|U_{M}|-|U_{0}|\right)$ edges.
Let us couple the inhomogeneous random graph $G$ with an Erdős-Rényi random graph $\widetilde{G}$ on $n$ nodes with marginal distribution $\widetilde{G}\sim\text{ER}(n,p_{\max})$, such that $G$ is a subgraph of $\widetilde{G}$. Thus,
(($\star$)( ⋆ ) ( $\leq\mathbb{P}\left(\exists\text{ subgraph $H\!=\!(W,F)$ of }G-v\text{ s.t. }|W|=3n^{\delta^{\prime}}\text{ and }|F|\geq 3\left(3n^{\delta^{\prime}}-|U_{0}|\right)\right)$
$\leq\mathbb{P}\left(|U_{0}|>n^{\delta^{\prime}}\right)+\mathbb{P}\left(\exists\text{ subgraph $H=(W,F)$ of }G\text{ s.t. }|W|=3n^{\delta^{\prime}}\text{ and }|F|\geq 6n^{\delta^{\prime}}\right)$
$\leq\mathbb{P}\left(|U_{0}|>n^{\delta^{\prime}}\right)+\mathbb{P}\left(\exists\text{ subgraph $\widetilde{H}\!=\!(\widetilde{W},\widetilde{F})$of }\widetilde{G}\text{ s.t. }|\widetilde{W}|=3n^{\delta^{\prime}}\text{ and }|\widetilde{F}|\geq 6n^{\delta^{\prime}}\right)$
$\leq o(1/n)+\underbrace{\binom{n}{3n^{\delta^{\prime}}}\cdot\mathbb{P}\left(\text{Bin}\left(\binom{3n^{\delta^{\prime}}}{2},p_{\max}\right)>6n^{\delta^{\prime}}\right)}_{(\star\star)},$
where the last step uses Lemma 16 and a union bound over all possible choices of $W$. Finally, using that $\binom{n}{k}\leq\left(\frac{ne}{k}\right)^{k}$ and the fact that $\text{Bin}(n,p)\preceq\text{Bin}(n^{\prime},p^{\prime})$ whenever $n^{\prime}\geq n$ and $p^{\prime}\geq p$ yields
[…]
$=\left(\frac{3e}{4}\cdot n^{\alpha/2-\varepsilon-(1-\delta^{\prime})/2}\right)^{6n^{\delta^{\prime}}}=o(1/n),$
since $\alpha/2-\varepsilon-(1-\delta^{\prime})/2<0$. Note (a) uses the concentration bound (17) from Lemma 17. This concludes the proof of Lemma 19. $\Box$
Finally, Lemma 19 directly implies Lemma 18.
4.3 Relating high degree nodes to the $k$-core: Proof of Lemma 13
Proof of Lemma 13. For any set $A\subseteq[n]$, let $N_{v}(A)$ denote the number of neighbors of $v$ in $A$, i.e.
$N_{v}(A):=|\left\{u\in A:\{u,v\}\in E(G)\right\}|.$
Fix $v\in[n]$. Let $G-v$ denote the induced subgraph of $G$ on the vertex set $[n]-\{v\}$. Let $C_{v}$ denote the $k$-core of $G-v$. Since $C_{v}$ is a subset of the $k$-core of $G$, it follows that
$\left\{v\notin\mathrm{core}_{k}(G)\right\}\subseteq\left\{N_{v}(\mathrm{core}_{k}(G))\leq k-1\right\}\subseteq\left\{N_{v}(C_{v})\leq k-1\right\}.$
Let $\delta\in(1-\alpha,1-\alpha+2\varepsilon)$. It follows that
($\mathbb{P}\left(\left\{v\notin\mathrm{core}_{k}(G)\right\}\cap\left\{\delta_{G}(v)\geq k+2/\alpha\right\}\right)\leq q_{1}+q_{2},$ (20)
where
$q_{1}:=\mathbb{P}\left(|C_{v}|<n-3n^{\delta}\right)\stackrel{{\scriptstyle\text{(a)}}}{{=}}o(1/n),$
where (a) follows from Lemma 18, and
$q_{2}$ :=($:=\mathbb{P}\left(\left\{N_{v}(C_{v})\leq k-1\right\}\cap\left\{\delta_{G}(v)\geq k+2/\alpha\right\}\cap\left\{|C_{v}|\geq n-3n^{1-\alpha/2}\right\}\right)$
$\leq\sum_{\begin{subarray}{c}A\subseteq V\\ |A|\geq n-3n^{\delta}\end{subarray}}\mathbb{P}\left(\left\{N_{v}(C_{v})<k\right\}\cap\left\{\delta_{G}(v)\geq k+2/\alpha\right\}\mid C_{v}=A\right)\cdot\mathbb{P}\left(C_{v}=A\right)$
$\leq\max_{\begin{subarray}{c}A\subseteq V\\ |A|\geq n-3n^{\delta}\end{subarray}}\mathbb{P}\left(\left\{N_{v}(A)<k\right\}\cap\left\{\delta_{G}(v)\geq k+2/\alpha\right\}\right)$
$\leq\max_{\begin{subarray}{c}A\subseteq V\\ |A|\geq n-3n^{\delta}\end{subarray}}\mathbb{P}\left(N_{v}(A^{c})\geq 1+2/\alpha\right)$ (21)
For all sets $A$ of size larger than $n-3n^{\delta}$, we have that
$\mathbb{P}\left(N_{v}(A^{c})\geq 1+2/\alpha\right)\!=\!\mathbb{P}\left(\sum_{i\in A^{c}}G_{vi}\geq 1+2/\alpha\right)\!\leq\mathbb{P}\left(\text{Bin}(3n^{\delta},n^{\alpha/2-1})\geq 1+2/\alpha\right),$ (22)
where the last step uses that $\sum_{i\in A^{c}}G_{vi}\preceq\text{Bin}(|A^{c}|,p_{\max})\preceq\text{Bin}(3n^{\delta},n^{\alpha/2-1})$. Combining (21) and (22) yields
$q_{2}$ $\leq\mathbb{P}\left(\text{Bin}(3n^{\delta},n^{\alpha/2-1})\geq 1+2/\alpha\right)\leq\left(\frac{3\alpha e}{2+\alpha}\cdot n^{-\alpha/2}\right)^{1+2/\alpha}\!=\text{const.}\times n^{-1-\alpha/2}=o(1/n).$
[…]
5 Proofs for Main Results
This section presents proofs for Theorems 1 and 8. Section 5.1 presents the intuition behind the analysis of Algorithm 1, collecting useful lemmas along the way. These lemmas are proved in Section 5.3, after the proof of Theorem 1 in Section 5.2. Finally, the negative result (Theorem 8) is proved in Section 5.5.
5.1 Analyzing Algorithm 1: Intuition behind the proof of Theorem 1